When the price of good A is P 40/ pair, the quantity for good B is 500 units. When the price increases by P 10 per unit, the demand for good B increases by 100 units. Compute the cross-price elasticity and determine the type of good.


1
Expert's answer
2022-02-02T14:43:21-0500

"E=\frac{\frac{600}{500}}{\frac{50}{40}}=0.96"


goods are called interchangeable

Е>0